# Social Bias in Popular Question-Answering Benchmarks

### v2

@@ 4.1 Benchmark Paper Analysis Results
We obtained two sets of annotations for each of the 30 benchmark papers, one by an internal annotator (first author) and one by an external annotator (inter-annotator agreement: $\kappa$=.78; $SD$=.10). [^13] Cohen’s $\kappa$ was computed on the basis of all yes-no questions excluding the ”suggest other annotation” category.Throughout this section, we present the internal annotations unless otherwise specified and only discuss some of the differences between internal and external annotations (all external results are presented in Appendix B).
4.1.1 Benchmark Motivation
Increased task difficulty, novelty, and more realistic problems were the most frequent reported motivations behind the benchmarks. Other motivating factors mentioned were increased benchmark dataset size, explainability/interpretability, and domain-specificity. We found none of the benchmark papers to be motivated by better social representativeness. Note that the external annotator found SIQA to be aiming for social representativeness since it is framed as a social intelligence benchmark (Sap et al., 2019). However, we did not find any evidence that the intention was to improve representativeness in a demographic sense.
@@ 4.1.2 Benchmark Creation and Annotation
From the 30 analyzed benchmark papers, 20 of the benchmarks consist of human-authored items. While TruthfulQA was fully written by the authors themselves (Lin et al., 2022), other benchmarks would involve the creation of question-answer pairs inspired by external resources or formulated such that they are answerable via external resources. In 13 cases, some type of web source was used as a basis. Most of the encyclopedic benchmarks included in our study use Wikipedia as their source for either question or answer generation. SQuAD v1.1 (Rajpurkar et al., 2016) consists of more than 100,000 questions about Wikipedia articles, posed by crowdworkers. Similarly, for StrategyQA (Geva et al., 2021), HotpotQA (Yang et al., 2018), and TruthfulQA (Lin et al., 2022), crowdworkers created question-answer pairs inspired by Wikipedia content. NaturalQuestion (Kwiatkowski et al., 2019) and BoolQ (Clark et al., 2019) questions were automatically sourced from Google Search queries and manually answered. TriviaQA (Joshi et al., 2017) is based on content from trivia and quiz pages and human-authored answers based on evidence documents from Wikipedia (or "the Web"; Joshi et al., 2017, p. 1602). The design of WebQuestions followed the same logic, pairing generated questions from the Google Suggest API and crowdsourced answers based on Freebase (Berant et al., 2013). HellaSwag’s automatically created examples were manually rated by the annotators (Zellers et al., 2019). Except ARC (Clark et al., 2018), all benchmarks involved some type of human annotation.
4.1.3 Benchmark Language
All but one of the selected benchmarks were in English, only. However, only 12 of the benchmark papers explicitly state this information. In all other cases, we had to derive this information from data examples. For these cases, we have to assume that the recruited annotators were sufficiently capable of understanding and following English instructions and writing and labeling English data examples. An exception to English as a default, is XQuAD, a multilingual benchmark based on translations of the English SQuAD v1.1 (Spanish, German, Greek, Russian, Turkish, Arabic, Vietnamese, Thai, Chinese, and Hindi; Artetxe et al., 2020). Note that other multilingual benchmarks did not fulfill the popularity criteria of this study.
@@ 4.1.4 Annotator Recruitment Criteria
For 50% of the benchmarks, crowdworkers were hired through Amazon Mechanical Turk. [^14] https://www.mturk.com/ Other platforms used are Surge AI [^15] https://www.surgehq.ai/ (Cobbe et al., 2021) and Upwork [^16] https://www.upwork.com/ (Rein et al., 2023). Again, only 15 benchmark papers mention criteria for the selection of annotators (see Table 1). These would include performance on the task, e.g., appraised in a screening test (Reddy et al., 2019), or their ratings on the crowdworking platform (Rein et al., 2023). Sometimes annotators were recruited due to their availability as co-authors or colleagues (Gordon et al., 2012; Lin et al., 2022; Yue et al., 2024). Another reason for recruitment would be expertise in a certain domain. BioASQ-QA, for example, is a biomedical benchmark that is fully written by domain experts (Krithara et al., 2023). It is reported where and in what type of institutions the experts hold positions (European universities, hospitals, and research institutes) as well as their concrete areas of research (e.g., "cardiovascular endocrinology, psychiatry, psychophysiology, pharmacology", p. 3). In StrategyQA, the authors refer to themselves as expert annotators (Geva et al., 2021). In other instances, what defines an expert is less clear. For example, in the OpenBookQA benchmark paper it is stated that the data were "filtered by an in-house expert to ensure higher quality" (Mihaylov et al., 2018, p. 2384) without further elaboration.
@@ 4.1.5 Annotator Demographics
Out of the 29 benchmark papers involving human annotators, 17 failed to report any demographic information (see Table 1). Country of recruitment or origin was mentioned for SQuAD, DROP, OpenBookQA, and MATH, exclusively referring to the USA, Canada, or North America in general (Rajpurkar et al., 2016; Dua et al., 2019; Mihaylov et al., 2018; Hendrycks et al., 2021). Level of education was mentioned in OpenBookQA (Master’s; Mihaylov et al., 2018), GPQA (PhD or higher; Rein et al., 2023), and MMMU (college students; Yue et al., 2024), which are based on textbook problems or exam knowledge. Information on age, gender, and ethnicity were not identified in the benchmark papers (by internal nor external annotator). Another indicator for demographic aspects are the author affiliations. We found that those are centered around renown North-American research institutes, universities, and technology firms. In sum, 13 of the benchmark papers were co-authored by researchers affiliated to the Allen Institute for Artificial Intelligence (Allen AI) and 8 by researchers affiliated to the University of Washington (UW).
@@ 4.1.6 Benchmark Bias and Toxicity
We asked annotators to answer the following question and include evidence for their answer: "Are analyses of aspects related to social bias, representativeness or toxicity in the benchmark dataset reported and, if so, what type of analyses?" The external annotators identified 4 benchmarks as informative in this regard. However, we noticed that they appeared to work on a different understanding of bias than us. For instance, OK-VQA utilizes (non-specific) label balancing to avoid heuristic prediction behavior [^17] For example, the question ”What season is it?” was mostly accompanied by the answer ”Winter” incentivizing the model to default to this answer (Marino et al., 2019). and for NaturalQuestions, an in-depth analyses of annotation variability was conducted. This indeed can be done in a social bias-sensitive manner (Haliburton et al., 2024), but in this case the focus was on general annotation quality (Kwiatkowski et al., 2019). We, hence, count these cases as uninformative of social bias or toxicity aspects.
We finally identified 3 out of 30 benchmark papers that clearly flag social biases in their data. [^18] None of the benchmark papers mentioned any toxicity-related metric (full agreement between internal and external annotations). The WinoGender bias metric (Rudinger et al., 2018) was applied to models trained on the WinoGrande train split (Sakaguchi et al., 2021) to verify its relative gender-fairness. The QuAC datasheet mentions potential biases towards famous men in its dataset as well as other not further specified biases.191919quac.ai/datasheet.pdf The GPQA benchmark paper explicitly states that bias was not avoided during the dataset creation. The authors "make no claim that GPQA is a representative sample of any population of questions that are likely to come up in the course of scientific practice," (Rein et al., 2023, p. 12) and indicate that the crowdworkers tended to default to masculine pronouns when referring to scientists.
An additional keyword matching for the terms "diverse" and "diversity" yielded matches in two thirds of the benchmark papers: Several pay attention to domain or topic diversity (e.g., Geva et al., 2021; Lu et al., 2022; Lin et al., 2022), question or answer diversity (e.g., Zellers et al., 2019; Bisk et al., 2019; Artetxe et al., 2020), as well as lexical diversity (e.g., Reddy et al., 2019; Dua et al., 2019; Cobbe et al., 2021). Yet, again, none of them account for demographic diversity.
@@ 4.2 Benchmark Data Analysis Results
Next, we analyzed distributions of gender, occupation, religion and location properties found for entities across 20 benchmark datasets (see Table 2, Appendix A), following the procedure described in Section 3.2.202020The selection of demographic markers reflects dimensions that are frequently discussed in the social bias in NLP literature (Sheng et al., 2021). The absolute number of entities differs greatly between benchmarks (see Table 6, Appendix B) due to differences in dataset sizes or the nature of the contents, e.g., HotpotQA (>20k entities) is inherently related to Wikipedia and, thus, highly overlaps with Wikidata, but MMLU (<100 entities) is designed to include knowledge not found on Wikipedia [^21] For MMLU, we only matched entities of type class of anatomical entity, cell type, and the like. and the commonsense benchmark COPA (<100 entities) does mostly not rely on real-world entities in its examples. [^22] Example: ”The man dropped on the floor. What happened as a result? For other commonsense and scholarly benchmarks we were in fact able to retrieve large numbers of entities suitable for our analysis: E.g., 893 human and 365 fictional human entities in SIQA, 480 human and 99 fictional human instances in GSM8K, or 317 mentions of U.S. States in ScienceQA.
@@
4.2.1 Gender and Occupation
Figure 3 shows the male-versus-female gender proportions across benchmarks. We only included benchmark datasets for which we found more than 30 gender entries. Genders beyond the binary were none or close to none and, hence, also not illustrated in the plot. The most favorable gender ratios are found in WinoGrande and HellaSwag, which are both commonsense benchmarks. As discussed in Section 4.1.6, a low gender bias metric was reported by the WinoGrande authors and our results confirm this observation. All Wikipedia-based benchmarks, like DROP, SQuAD, or TriviaQA exhibit prominent gender gaps. In fact, the DROP benchmark dataset is only based on text passages about male-dominated "National Football League (NFL) game summaries and history articles" (Dua et al., 2019, p. 2371).
@@ 4.2.1 Gender and Occupation
Additionally, we found that the most represented occupations differ for female and male entities. For example for GSM8K (school math problems) and WinoGrande (commonsense), the male top-10 occupations for both benchmarks include several athletic professions, while the female top-10 occupations are biased towards entertainment roles. Visualizations of occupation-related gender biases for eight different benchmarks are provided in Figure 5, Appendix C.
4.2.2 Religion and Location
As indicators of cultural and geographic context, we examined religions and locations. Here, we examined the benchmarks for which more than 30 religion properties were retrieved (ranging between 33 for BoolQ and 652 for TriviaQA). Christianity and instances of Christian religions rank highest across benchmarks. In fact, Christianity and/or Catholicism are among the top-3 religion labels for 14 out of 15 benchmarks. All distributions are visualized in Figure 6, Appendix C.
Across encyclopedic, commonsense, and scholarly benchmarks, most coordinates are located around North America and Western Europe and Eastern and Southern regions are less represented. For HotpotQA, TriviaQA, and NaturalQuestions slightly more coordinates are located on the South American, African, and Australian continents compared to the other benchmarks. Map visualizations for 17 benchmarks are provided in Figure 7, Appendix C. We also retrieved location names associated to entities in the datasets. Again, Western regions are more represented. E.g., for BoolQ and StrategyQA, the most frequently named location are the United States (56% and 31%) and the United Kingdom (9% and 15%), followed by Canada (2%) for BoolQ and Brazil, and Japan (4% each) for StrategyQA.
@@ 5 Discussion
Our study reveals significant gaps in transparency as well as harmful social biases across the 30 most popular QA and RC benchmarks. While two-thirds consist of human-authored examples and nearly all involve human annotation, half of them fail to disclose annotator demographics or recruitment criteria and all fail to report gender, age, or ethnicity. For the remaining cases, countries of origin or recruitment are predominantly North American, reflecting the Western institutional affiliations of the benchmark authors. While the QuAC paper stands out for its thorough reporting, others like MMLU (which is commonly referenced to market flagship models of famous tech firms [^23] https://openai.com/index/hello-gpt-4o/ [^24] https://www.anthropic.com/news/3-5-models-and-computer-use) lack all of the details we were looking for. The benchmarks overwhelmingly prioritize task difficulty or novelty aspects over social representativeness. The reliance on Wikipedia (known for its representational issues; Sun and Peng, 2021; Menking and Rosenberg, 2021; Tripodi, 2023) for encyclopedic benchmarks, perpetuates underrepresentation of marginalized communities and introduces biases in gender, occupation, geography, and religion. But also commonsense and scholarly benchmarks were found to default to male and Western examples. All but one benchmark are fully English, but less than half report this information in the paper, disregarding that they are indeed language-specific benchmarks (Bender, 2011).
Our findings highlight a systemic lack of diversity and transparency in widely used QA and RC benchmarks. These shortcomings perpetuate the development of technologies that produce harmful, discriminatory outcomes. Furthermore, our findings exemplify once more a "laissez-faire attitude" (Paullada et al., 2021, p. 4) prevalent in AI dataset creation, which must be addressed through the implementation of robust documentation, validation, and representation standards. While we acknowledge the growing discourse around better AI evaluation practices (Wallach et al., 2024), we emphasize that the conversation must prioritize social bias alongside validity and transparency.

### v3

@@ 4.1 Benchmark Paper Analysis Results
We obtained two sets of annotations for each of the 30 benchmark papers, one by an internal annotator (first author) and one by an external annotator (inter-annotator agreement: $\kappa$=.78; SDSD=.10). [^15] Cohen’s $\kappa$ was computed on the basis of all yes-no questions excluding the ”suggest other annotation” category.
Throughout this section, we present the internal annotations unless otherwise specified and only discuss some of the differences between internal and external annotations (all external results are presented in Appendix B).
@@ 4.1.1 Benchmark Creation and Annotation
To answer RQ1, we firstly examined how the benchmark data and annotations were sourced. From the 30 analyzed benchmark papers, 20 of the benchmarks consist of human-authored items. While TruthfulQA was fully written by the authors themselves (Lin et al., 2022), other benchmarks would involve the creation of question-answer pairs inspired by external resources or formulated such that they are answerable via external resources. In 13 cases, some type of web source was used as a basis. Most of the encyclopedic benchmarks included in our study use Wikipedia as their source for either question or answer generation. SQuAD v1.1 (Rajpurkar et al., 2016) consists of more than 100,000 questions about Wikipedia articles, posed by crowdworkers. Similarly, for StrategyQA (Geva et al., 2021), HotpotQA (Yang et al., 2018), and TruthfulQA (Lin et al., 2022), crowdworkers created question-answer pairs inspired by Wikipedia content. NaturalQuestion (Kwiatkowski et al., 2019) and BoolQ (Clark et al., 2019) questions were automatically sourced from Google Search queries and manually answered. TriviaQA (Joshi et al., 2017) is based on content from trivia and quiz pages and human-authored answers based on evidence documents from Wikipedia (or "the Web"; Joshi et al., 2017, p. 1602). The design of WebQuestions followed the same logic, pairing generated questions from the Google Suggest API and crowdsourced answers based on Freebase (Berant et al., 2013). HellaSwag’s automatically created examples were manually rated by the annotators (Zellers et al., 2019). Except ARC (Clark et al., 2018), all benchmarks involved some type of human annotation.
@@ 4.1.2 Annotator Recruitment Criteria
Another important factor to consider with respect to RQ1 are the criteria by which annotators were recruited. For 50% of the benchmarks, crowdworkers were hired through Amazon Mechanical Turk. [^16] https://www.mturk.com/ Other platforms used are Surge AI [^17] https://www.surgehq.ai/ (Cobbe et al., 2021) and Upwork [^18] https://www.upwork.com/ (Rein et al., 2023). Again, only 15 benchmark papers mention criteria for the selection of annotators (see Table 1). These would include performance on the task, e.g., appraised in a screening test (Reddy et al., 2019), or their ratings on the crowdworking platform (Rein et al., 2023). Sometimes annotators were recruited due to their availability as co-authors or colleagues (Gordon et al., 2012; Lin et al., 2022; Yue et al., 2024). Another reason for recruitment would be expertise in a certain domain. BioASQ-QA, for example, is a biomedical benchmark that is fully written by domain experts (Krithara et al., 2023). It is reported where and in what type of institutions the experts hold positions (European universities, hospitals, and research institutes) as well as their concrete areas of research (e.g., "cardiovascular endocrinology, psychiatry, psychophysiology, pharmacology", p. 3). In StrategyQA, the authors refer to themselves as expert annotators (Geva et al., 2021). In other instances, what defines an expert is less clear. For example, in the OpenBookQA benchmark paper it is stated that the data were "filtered by an in-house expert to ensure higher quality" (Mihaylov et al., 2018, p. 2384) without further elaboration.
@@ 4.1.3 Annotator Demographics
Finally, we looked for potentially reported demographic details to learn more about the identity or situatedness of those involved in the benchmark creation (relevant for RQ1, as well as RQ3). Out of the 29 benchmark papers involving human annotators, 17 failed to report any demographic information (see Table 1). Country of recruitment or origin was mentioned for SQuAD, DROP, OpenBookQA, and MATH, exclusively referring to the USA, Canada, or North America in general (Rajpurkar et al., 2016; Dua et al., 2019; Mihaylov et al., 2018; Hendrycks et al., 2021). Level of education was mentioned in OpenBookQA (Master’s; Mihaylov et al., 2018), GPQA (PhD or higher; Rein et al., 2023), and MMMU (college students; Yue et al., 2024), which are based on textbook problems or exam knowledge. Information on age, gender, and ethnicity were not identified in the benchmark papers (by internal nor external annotator). Another indicator for demographic aspects are the author affiliations. We found that those are centered around renown North-American research institutes, universities, and technology firms. In sum, 13 of the benchmark papers were co-authored by researchers affiliated to the Allen Institute for Artificial Intelligence (Allen AI) and 8 by researchers affiliated to the University of Washington (UW).
4.1.4 Benchmark Motivation
With reference to RQ2, we were interested to learn what motivated creators to develop their specific benchmarks, and whether or not any of the benchmarks was motivated by an aim to achieve good social representativeness. This was not found to be the case for any of the benchmark papers. Note that the external annotator found SIQA to be aiming for social representativeness since it is framed as a social intelligence benchmark (Sap et al., 2019). However, we did not find any evidence that the intention was to improve representativeness in a demographic sense.
Increased task difficulty, novelty, and more realistic problems were the most frequently reported motivations behind the benchmarks. Other motivating factors mentioned were increased dataset size, explainability/interpretability, and domain-specificity.
@@ 4.1.5 Benchmark Bias and Toxicity
In reference to RQ2, we also asked annotators to answer the following question and include evidence for their answer: "Are analyses of aspects related to social bias, representativeness or toxicity in the benchmark dataset reported and, if so, what type of analyses?" The external annotators identified 4 benchmarks as informative in this regard. However, we noticed that they appeared to work on a different understanding of bias than us. For instance, OK-VQA utilizes (non-specific) label balancing to avoid heuristic prediction behavior [^19] For example, the question ”What season is it?” was mostly accompanied by the answer ”Winter” incentivizing the model to default to this answer (Marino et al., 2019). and for NaturalQuestions, an in-depth analyses of annotation variability was conducted. This indeed can be done in a social bias-sensitive manner (Haliburton et al., 2025), but in this case the focus was on general annotation quality (Kwiatkowski et al., 2019). We count these as uninformative of social bias or toxicity aspects.
We finally identified 3 out of 30 benchmark papers that clearly flag social biases in their data. [^20] None of the benchmark papers mentioned any toxicity-related metric (full agreement between internal and external annotations). The WinoGender bias metric (Rudinger et al., 2018) was applied to models trained on the WinoGrande train split (Sakaguchi et al., 2021) to verify its relative gender-fairness. The QuAC datasheet mentions potential biases towards famous men in its dataset as well as other not further specified biases.212121quac.ai/datasheet.pdf The GPQA benchmark paper explicitly states that bias was not avoided during the dataset creation. The authors "make no claim that GPQA is a representative sample of any population of questions that are likely to come up in the course of scientific practice," (Rein et al., 2023, p. 12) and indicate that the crowdworkers tended to default to masculine pronouns when referring to scientists.
An additional keyword matching for the terms "diverse" and "diversity" yielded matches in two thirds of the benchmark papers: Several pay attention to domain or topic diversity (e.g., Geva et al., 2021; Lu et al., 2022; Lin et al., 2022), question or answer diversity (e.g., Zellers et al., 2019; Bisk et al., 2020; Artetxe et al., 2020), as well as lexical diversity (e.g., Reddy et al., 2019; Dua et al., 2019; Cobbe et al., 2021). Yet, again, none of them account for demographic diversity.
4.1.6 Benchmark Language
Finally, the language of the benchmark is another factor that can be indicative of a socially relevant form of bias, namely linguistic bias (RQ2). All but one of the selected benchmarks were in English, only. Yet, only 12 of the benchmark papers explicitly state this information. In all other cases, we had to derive this information from data examples. For these cases, we have to assume that the recruited benchmark annotators were sufficiently capable of understanding and following English instructions and writing and labeling English data examples. An exception to English as a default, is XQuAD, a multilingual benchmark based on translations of the English SQuAD v1.1 (Spanish, German, Greek, Russian, Turkish, Arabic, Vietnamese, Thai, Chinese, and Hindi; Artetxe et al., 2020). Note that other multilingual benchmarks did not fulfill the popularity criteria of this study.
@@ 4.2 Benchmark Data Analysis Results
To find more evidence towards answering RQ2 and RQ3, we analyzed distributions of gender, occupation, religion and location properties found for entities across 20 benchmark datasets (see Table 2, Appendix A), following the procedure described in Section 3.2.222222The selection of demographic markers reflects dimensions that are frequently discussed in the social bias in NLP literature (Sheng et al., 2021). The absolute number of entities differs greatly between benchmarks (see Table 6, Appendix B) due to differences in dataset sizes or the nature of the contents, e.g., HotpotQA (>20k entities) is inherently related to Wikipedia and, thus, highly overlaps with Wikidata, but the commonsense benchmark COPA (<100 entities) does mostly not rely on real-world entities in its examples. [^23] Example: ”The man dropped on the floor. What happened as a result? [^24] The entity linker was validated on a small subset of data, consisting of 50 randomly selected items from each benchmark. The first author annotated these random samples by manually listing the Wikidata QIDs associated to entities mentioned therein. The Micro F1 score across benchmarks is .73 (precision=.63, recall=.87).
@@
4.2.1 Gender
Figure 3 shows the male-to-female gender ratios across benchmarks. We only included benchmark datasets for which we found more than 30 gender entries. Genders beyond the binary were none or close to none and not illustrated in the plot. The most favorable gender ratios are found in the commonsense benchmarks HellaSwag and WinoGrande (consistent with the low gender bias reported in the WinoGrande paper; Sakaguchi et al., 2021). All Wikipedia-based benchmarks, like DROP, SQuAD, or TriviaQA exhibit prominent gender gaps. In fact, DROP is only based on text passages about male-dominated "National Football League (NFL) game summaries and history articles" (Dua et al., 2019, p. 2371).
@@ 4.2.1 Gender
Additionally, we found that the most frequent occupations differ for female and male entities. For example, for WinoGrande (commonsense), the top-10 male occupations include several athletic professions, while the top-10 female occupations are leaning towards entertainment roles (see Figure 5, Appendix C).
4.2.2 Religion
As an indicator of cultural context, we examined the distributions of religions. Firstly, we determined the benchmarks for which 30 or more religion properties were retrieved (ranging between 33 for BoolQ and 652 for TriviaQA). Christianity and instances of Christian religions rank highest across benchmarks. In fact, Christianity and/or Catholicism are among the top-3 religion labels for 14 out of the 15 benchmarks (see Figure 6, Appendix C). Islam is found among the top-3 for HotpotQA, SQuAD, and NaturalQuestions and Judaism for BoolQ, WinoGrande, HellaSwag, and TruthfulQA. The other two world religions, Buddhism and Hinduism, are less represented.
4.3 Location
For the analysis of locations, we again filtered for benchmarks with at least 30 matched location properties. Across encyclopedic, commonsense, and scholarly benchmarks, most coordinates are located around North America and Western Europe. Eastern and Southern regions are less represented. For HotpotQA, TriviaQA, and NaturalQuestions slightly more coordinates are located on the South American, African, and Australian continents compared to the other benchmarks (see Figure 7, Appendix C). We also retrieved location names associated to entities in the datasets. Again, Western regions are more represented. E.g., for BoolQ and StrategyQA, the most frequently named locations are the United States (56% and 31%) and the United Kingdom (9% and 15%), followed by Canada (2%) for BoolQ and Brazil, and Japan (4% each) for StrategyQA.
@@ 5 Discussion
(RQ1) Most of the benchmarks consist of human-authored examples and nearly all involve human annotation. Yet, demographic and recruitment details are rarely reported. While the QuAC paper stands out for its comprehensive reporting, several others like MMLU (which is commonly referenced to market flagship models of famous tech firms) [^25] https://openai.com/index/hello-gpt-4o/ [^26] https://www.anthropic.com/news/3-5-models-and-computer-use lack all of the details we were looking for. In a few cases, countries of origin/recruitment are reported (mostly North American). These observations emphasize once more that benchmark creators are not sufficiently sensitized towards the situatedness of their practice (Raji et al., 2021). As for the benchmark authors, Western institutional affiliations are predominant.
(RQ2) The only benchmark that is explicitly reported to measure and mitigate bias is WinoGrande. It utilizes the WinoGender metric to control for (binary) gender bias. Several of the remaining benchmarks datasets are biased regarding gender, occupation, religion, and location of the entities of interest. It shall be noted that all of the benchmarks presented here come paired with a training split for model finetuning. Hence, the biases affect not only evaluation but also training. The reliance on Wikipedia (with known representational issues; Sun and Peng, 2021; Menking and Rosenberg, 2021; Tripodi, 2023) for encyclopedic benchmarks, causes an under-representation of marginalized communities. But also commonsense and scholarly benchmarks were found to default to male and Western examples.
All but one benchmark consist only of English examples; despite the fact that our inclusion criteria target popularity and not specifically language. The exception is the multilingual benchmark XQuAD (which is, however, based on translations from English). Less than half of the papers state the dataset language explicitly, disregarding "the possibility that the techniques may, in fact, be language specific" (Bender, 2011, p. 18). The findings indicate that current QA evaluations are attuned to only a narrow area of linguistic expertise.
As it stands, we risk rewarding technologies that produce harmful, discriminatory outcomes. Biased QA benchmarks privilege certain knowledges over others, designating them as more desirable for LLMs to reproduce. Such LLMs (e.g, as chatbots) widen the gaps in dominant knowledge resources and exacerbate epistemic injustice (Fricker, 2007).
(RQ3) Previous studies have demonstrated the influence of annotator demographics on annotations (Sap et al., 2022; Pei and Jurgens, 2023; Al Kuwatly et al., 2020). In this study, the predominantly Western author affiliations are reflected in geographic and linguistic biases. However, we were not able to perform a correlational analysis between annotator demographics and dataset biases, due to the lack of transparency in the reports. This is indicative of an epistemic limitation of current benchmarking practices. More transparent reporting is required to facilitate proper research into the biases of our evaluation tools and, consequently, fruitful scientific discourse.
Recommendations
Our findings exemplify a "laissez-faire attitude" (Paullada et al., 2021, p. 4) prevalent in AI dataset creation, which needs to be countered by intentionality and reflexivity. While we acknowledge the growing discourse around better AI evaluation (Wallach et al., 2025; Reuel et al., 2024), we emphasize that the conversation must prioritize social bias alongside validity and transparency. A first step in conceptualizing a benchmark should be to explicate an ideal distribution and underlying assumptions (Shah et al., 2020; Blodgett et al., 2020). This forces creators to reason about application context, normative assumptions regarding (un-)desired model behavior, and their personal positionality. Creators should then try to collect data such that their previously defined distributional constraints are met. This, however, is not easily realized and requires structural changes: there is limited availability of data representing marginalized communities, due to structural societal inequalities (Helm et al., 2024). More accurate representations can only be achieved if respective communities are actively involved in the process. This must be realized through non-exploitative, true participation (Birhane et al., 2022a). We argue that there is not only ethical but also epistemic value in pursuing respective efforts, as this helps to foster more representative and generalizable evaluation (Harding, 1986). Limitations and biases are always expected. Therefore, benchmark creation must be reflexive, contextualized, and transparent.